If a boy tosses a ball straight up with an initial velocity of 13m/s what .how long does the ball take at it's highest


1
Expert's answer
2021-02-17T10:41:30-0500

The velocity of the ball before it reaches the highest point is given as follows:


"v = v_0 -gt"

where "v_0 = 13m\/s" is the initial speed, "g = 9.81m\/s^2" is the gravitational acceleration, and "t" is the time. At the highest point "v = 0". Thus, the moment of time is:


"0 = v_0 - gt_h\\
t_h = \dfrac{v_0}{g}\\
t_h = \dfrac{13}{9.81}\approx 1.33s\\"

Answer. 1.33s